# Large Language Models Enhanced Hyperbolic Space Recommender Systems

### v1

@@ Abstract.
Large Language Models (LLMs) have attracted significant attention in recommender systems for their excellent world knowledge capabilities. However, existing methods that rely on Euclidean space struggle to capture the rich hierarchical information inherent in textual and semantic data, which is essential for capturing user preferences. The geometric properties of hyperbolic space offer a promising solution to address this issue. Nevertheless, integrating LLMs-based methods with hyperbolic space to effectively extract and incorporate diverse hierarchical information is non-trivial. To this end, we propose a model-agnostic framework, named HyperLLM, which extracts and integrates hierarchical information from both structural and semantic perspectives. Structurally, HyperLLM uses LLMs to generate multi-level classification tags with hierarchical parent-child relationships for each item. Then, tag–item and user–item interactions are jointly learned and aligned through contrastive learning, thereby providing the model with clear hierarchical information. Semantically, HyperLLM introduces a novel meta-optimized strategy to extract hierarchical information from semantic embeddings and bridge the gap between the semantic and collaborative spaces for seamless integration. Extensive experiments show that HyperLLM significantly outperforms recommender systems based on hyperbolic space and LLMs, achieving performance improvements of over 40%. Furthermore, HyperLLM not only improves recommender performance but also enhances training stability, highlighting the critical role of hierarchical information in recommender systems.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13-18, 2025; Padua, Italyisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
From the structural perspective, we focus on well-defined hierarchical relationships in the data. Specifically, we use LLMs to generate multi-level tags with parent-child relationships for each item, representing classifications ranging from broad to specific. These tags are inspired by classification networks  (Horta et al., 2023) in real-world taxonomies, which exhibit clear hierarchical structures. To incorporate this information into the hyperbolic model, we construct a tag–item matrix and jointly learn the tag–item and user–item interactions. Subsequently, hyperbolic space contrastive learning  (Qin et al., 2024) aligns item representations from both sides, collectively providing structural hierarchical information to the model.
@@ 4.1. LLMs-based Structural Extraction
{mdframed}
[backgroundcolor=black!15] Item Prompt:
@@
{mdframed}
[backgroundcolor=black!15] User Prompt:
@@ 4.3. Structural and Semantic Integration
After the Meta-optimized Semantic Extraction module is completed, we integrate the extracted structural and semantic hierarchical information into the hyperbolic model, corresponding to the second phase of the meta-optimized strategy. In this phase, we use the MoE-transformed embeddings $\mathbf{x}_{u}$, $\mathbf{x}_{i}$ to initialize the user and item ID embeddings, and then tain the hyperbolic model.
@@ 5.1. Experimental Setup
Table 1. Dataset statistics. Avg Token denotes the average number of tokens in the text Features

### v2

@@ Abstract.
Large Language Models (LLMs) have attracted significant attention in recommender systems for their excellent world knowledge capabilities. However, existing methods that rely on Euclidean space struggle to capture the rich hierarchical information inherent in textual and semantic data, which is essential for capturing user preferences. The geometric properties of hyperbolic space offer a promising solution to address this issue. Nevertheless, integrating LLMs-based methods with hyperbolic space to effectively extract and incorporate diverse hierarchical information is non-trivial. To this end, we propose a model-agnostic framework, named HyperLLM, which extracts and integrates hierarchical information from both structural and semantic perspectives. Structurally, HyperLLM uses LLMs to generate multi-level classification tags with hierarchical parent-child relationships for each item. Then, tag-item and user-item interactions are jointly learned and aligned through contrastive learning, thereby providing the model with clear hierarchical information. Semantically, HyperLLM introduces a novel meta-optimized strategy to extract hierarchical information from semantic embeddings and bridge the gap between the semantic and collaborative spaces for seamless integration. Extensive experiments show that HyperLLM significantly outperforms recommender systems based on hyperbolic space and LLMs, achieving performance improvements of over 40%. Furthermore, HyperLLM not only improves recommender performance but also enhances training stability, highlighting the critical role of hierarchical information in recommender systems.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730019isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Recommender systems
@@ 1. Introduction
From the structural perspective, we focus on well-defined hierarchical relationships in the data. Specifically, we use LLMs to generate multi-level tags with parent-child relationships for each item, representing classifications ranging from broad to specific. These tags are inspired by classification networks  (Horta et al., 2023) in real-world taxonomies, which exhibit clear hierarchical structures. To incorporate this information into the hyperbolic model, we construct a tag-item matrix and jointly learn the tag-item and user-item interactions. Subsequently, hyperbolic space contrastive learning  (Qin et al., 2024) aligns item representations from both sides, collectively providing structural hierarchical information to the model.
@@ 4.1. LLMs-based Structural Extraction
Item Prompt:
@@
User Prompt:
@@ 4.3. Structural and Semantic Integration
After the Meta-optimized Semantic Extraction module is completed, we integrate the extracted structural and semantic hierarchical information into the hyperbolic model, corresponding to the second phase of the meta-optimized strategy. In this phase, we use the MoE-transformed embeddings $\mathbf{x}_{u}$, $\mathbf{x}_{i}$ to initialize the user and item ID embeddings, and then train the hyperbolic model.
@@ 5.1. Experimental Setup
Table 1. Dataset statistics. Avg Token denotes the average number of tokens in the text Features.